OpenStudy (anonymous):

Confused help thanks. Eva has 35 jelly beans in a bag. The bag has 12 lemon, 5 grape, 7 orange, 9 strawberry, and 2 coconut. If she chooses one jelly bean without looking, eats it, then chooses another jelly bean without looking, what is the probability that she will choose both coconut jelly beans?

OpenStudy (mertsj):

\[\frac{2}{35}\times\frac{1}{34}=\frac{1}{595}\]

OpenStudy (anonymous):

thank you

OpenStudy (mertsj):

yw